OpenStudy (anonymous):

Consider the given curves to do the following. x=4sqrt(y) x = 0 y = 1 Use the method of cylindrical shells to find the volume V of the solid obtained by rotating the region bounded by the given curves about the x-axis.

OpenStudy (anonymous):

Draw a diagram to visualize the problem.

OpenStudy (anonymous):

i put it so that its in x terms. so its y = (x^2)/16. am i able to do that?

OpenStudy (anonymous):

Yes you can

OpenStudy (anonymous):

\[2\pi \int\limits_{0}^{1}y(4y^(1/2))dy\]

OpenStudy (anonymous):

sorry its suppose to be 4y^1/2

OpenStudy (anonymous):

is that set up correctly?

OpenStudy (anonymous):

Let me calculate.

OpenStudy (anonymous):

Sorry, had to take a call. This has a little space between the function and x axis.

OpenStudy (anonymous):

its okay. and yeah it does

OpenStudy (anonymous):

How did you determine the boundary of 1?

OpenStudy (anonymous):

o, oops i think i got it confused for the y=0

OpenStudy (anonymous):

how do you determine the boundary for this integral?

OpenStudy (anonymous):

o wait sorry

OpenStudy (anonymous):

If given freedom you would use washer technique. Instructions call for cylindrical shells, I think, check on this\[\int\limits_{0}^{4}2\pi(1-4\sqrt{y})4\sqrt{y}\]